### v2

# Advancing Software Quality: A Standards-Focused Review of LLM-Based Assurance Techniques

Abstract
Software Quality Assurance (SQA) is critical for delivering reliable, secure, and efficient software products. The Software Quality Assurance Process aims to provide assurance that work products and processes comply with predefined provisions and plans. Recent advancements in Large Language Models (LLMs) present new opportunities to enhance existing SQA processes by automating tasks like requirement analysis, code review, test generation, and compliance checks. Simultaneously, established standards such as ISO/IEC 12207, ISO/IEC 25010, ISO/IEC 5055, ISO 9001/ISO/IEC 90003, CMMI, and TMM provide structured frameworks for ensuring robust quality practices. This paper surveys the intersection of LLM-based SQA methods and these recognized standards, highlighting how AI-driven solutions can augment traditional approaches while maintaining compliance and process maturity. We first review the foundational software quality standards and the technical fundamentals of LLMs in software engineering. Next, we explore various LLM-based SQA applications, including requirement validation, defect detection, test generation, and documentation maintenance. We then map these applications to key software quality frameworks, illustrating how LLMs can address specific requirements and metrics within each standard. Empirical case studies and open-source initiatives demonstrate the practical viability of these methods. At the same time, discussions on challenges (e.g., data privacy, model bias, explainability) underscore the need for deliberate governance and auditing. Finally, we propose future directions encompassing adaptive learning, privacy-focused deployments, multimodal analysis, and evolving standards for AI-driven software quality. By uniting insights from academic research, industry best practices, and established quality frameworks, we provide a comprehensive blueprint for integrating LLMs into SQA in a trustworthy, efficient, and standards-aligned manner.
I Introduction
Software systems continue to grow in complexity and scope, with modern applications often integrating numerous services and components [27]. As a result, Software Quality Assurance (SQA) activities must adapt to ensure reliability, security, and maintainability across distributed architectures. A key challenge arises from the increasing velocity of software releases, necessitating more automated and intelligent methods to maintain high-quality standards.
Large Language Models (LLMs) have emerged as a powerful asset in addressing these challenges. By leveraging advanced neural architectures—particularly the transformer model introduced by Vaswani et al. [143]—LLMs can learn contextual relationships from extensive natural language text and source code repositories. Recent works highlight their applicability to various software engineering tasks, including code completion and defect detection [24, 49]. For example, GitHub Copilot has demonstrated the potential to reduce coding effort by providing real-time suggestions, though questions remain regarding its accuracy and coverage [24].
The convergence of LLMs and SQA presents an opportunity to automate traditionally labor-intensive tasks, from identifying ambiguous requirements [111] to generating comprehensive test suites [139]. Although these approaches show promise, they must be systematically integrated into established frameworks and standards that guide quality assurance across industries [105]. Standards such as ISO/IEC 12207, ISO/IEC 25010, ISO/IEC 5055, ISO 9001 (and 90003), CMMI, and the Test Maturity Model (TMM) have long provided robust structures for ensuring software quality [79]. As organizations explore AI-driven methods, aligning these techniques with known best practices is crucial for consistent, reliable outcomes [69].
This survey examines how LLM-based techniques can enhance, extend, or streamline SQA activities within the context of recognized software quality standards. We begin with an overview of the major standards and then explore the fundamentals of large language models in software engineering. Subsequent sections delve into specific LLM-based approaches for requirements analysis, code review, testing, maintenance, and compliance. We present real-world case studies that illustrate the practical adoption of these methods, followed by an in-depth discussion of associated challenges, limitations, and risks. Finally, we conclude with prospective directions for research and policy, emphasizing the need for standards evolution to fully harness the potential of LLM-assisted SQA.
II Overview of Software Quality Standards
Software Quality Assurance (SQA) relies on established standards to guide the design, development, and maintenance of reliable, secure, and performant systems. This section provides an overview of key standards and models that underpin traditional SQA practices. In later sections, we will explore how large language model (LLM)- driven approaches can be aligned with or extended by these frameworks.
II-A ISO/IEC 12207 (Software Life Cycle Processes)
ISO/IEC 12207 was published on 1 August 1995. It was the first International Standard to provide a comprehensive set of life cycle processes, activities, and tasks for software that is part of a larger system and for stand-alone software products and services. ISO/IEC 12207 defines a comprehensive set of life cycle processes for software, spanning planning, development, operation, maintenance, and decommissioning [68]. It emphasizes a systematic and standardized approach throughout the software life cycle, mandating well-defined roles, responsibilities, and deliverables for each phase.  The standard underscores the importance of verification and validation (V&V) processes, ensuring that products meet user requirements and intended purposes [79]. By integrating SQA tasks into these life cycle processes, ISO/IEC 12207 provides a robust framework for achieving software quality from concept definition to retirement.
II-B ISO/IEC 25010 (Systems and Software Quality Requirements and Evaluation – SQuaRE)
Published under the SQuaRE series, ISO/IEC 25010 presents a quality model that classifies software product quality characteristics into eight main categories, including functional suitability, performance efficiency, compatibility, usability, reliability, security, maintainability, and portability [65].  The model aims to provide stakeholders with a shared understanding of essential quality attributes and how to measure them. Researchers have employed ISO/IEC 25010 as a basis for empirical evaluations of software products, revealing how different projects prioritize certain attributes (e.g., security over performance) depending on domain-specific needs [53].
II-C ISO/IEC 5055 (Measuring Internal Software Quality)
ISO/IEC 5055:2021 outlines a standard for detecting and measuring structural weaknesses in software systems by focusing on four critical factors: security, reliability, performance efficiency, and maintainability [66]. It formalizes metrics for issues such as control flow complexity and code duplication, providing quantitative thresholds for severity.  By systematically identifying code-level weaknesses, ISO/IEC 5055 serves as a critical tool for organizations looking to address deep-seated structural flaws early. In practice, combining static analysis with contextual insights—such as those gleaned from an LLM-based review system—can extend the utility of this standard beyond purely syntactic checks [104].
II-D ISO 9001 and ISO/IEC 90003
Although ISO 9001 applies to quality management systems across industries, its principles of defining processes, documenting procedures, and ensuring continuous improvement underpin software quality assurance [64]. ISO 9001 mandates risk-based thinking, leadership involvement, and clear documentation of quality objectives.  ISO/IEC 90003 interprets the requirements of ISO 9001 specifically for software development and maintenance [67]. It provides additional guidance on software-related processes, ensuring organizations align quality management systems with the unique challenges of building and evolving software.
II-E Capability Maturity Model Integration (CMMI)
Developed by the Software Engineering Institute (SEI), Capability Maturity Model Integration (CMMI) offers a structured approach for process improvement, grouping organizational practices into maturity levels [62]. Each level represents a higher degree of process capability, from ad hoc procedures (Level 1) to systematic optimization (Level 5).  CMMI emphasizes quantitative project management and continuous improvement. By integrating LLM-based analytics (e.g., defect prediction, requirement coverage) into CMMI-based practices, organizations can rapidly advance through maturity levels while maintaining robust oversight [38].
II-F Test Maturity Model (TMM)
The Test Maturity Model (TMM) is designed to assess and improve the maturity of testing processes within software organizations. It delineates stages of test process development, from initial and unstructured to fully optimized [21]. Each level specifies objectives and deliverables that elevate testing quality, such as test planning, monitoring, and control.  TMM’s emphasis on continuous assessment and structured improvements aligns with modern DevOps pipelines. Researchers have highlighted that automated tools, including AI-based test generation, can significantly enhance testing efficiency [53], suggesting that TMM-based organizations could benefit from integrating LLM-driven test automation and analysis.
III Large Language Models in Software Engineering
III-A Fundamentals of LLMs
Transformer Architectures.
The seminal work by Vaswani et al. [143] introduced the transformer model, a cornerstone of contemporary large language models (LLMs). Unlike recurrent neural networks, transformers leverage a self-attention mechanism to capture global dependencies in sequences, allowing models to process text (and code) in parallel rather than sequentially. This design significantly reduces training times and enhances the capacity to learn nuanced relationships from large-scale corpora.
Pre-training and Fine-tuning.
Modern LLMs undergo a two-stage process: pre-training on massive datasets, followed by fine-tuning for domain-specific tasks [35, 20]. Pre-training typically involves learning general linguistic or code-related representations while fine-tuning refines the model for specialized objectives such as code generation, bug detection, or requirement analysis [49].
Emerging Models and Frameworks.
Recent innovations have produced a variety of transformer-based models tailored for different applications:
  • GPT Series (OpenAI). In particular, GPT-3 [20] and GPT-4 focus on natural language generation, boasting billions of parameters trained on diverse internet-scale corpora.
  • BERT Variants (Google). BERT [35], and its successors (RoBERTa, DistilBERT) use bidirectional context analysis to excel in tasks like sentence classification and question answering.
  • CodeBERT, CodeT5, PolyCoder. Targeted at code-related tasks, CodeBERT [49] and CodeT5 [117] adapt transformer architectures to programming languages, enabling code completion and bug detection. PolyCoder [158] further explores model architectures optimized for lightweight code understanding and generation.
III-B Key LLM-based Tasks in Software Engineering
Code Generation and Completion.
GitHub Copilot [24] illustrates how advanced LLMs can assist developers by offering real-time suggestions. These suggestions can reduce coding effort, yet they also risk introducing subtle defects if developers become overly reliant on them [142]. Balancing productivity gains with thorough code review remains an open challenge.
Requirements Engineering and Analysis.
LLMs can parse large volumes of textual requirements, highlighting ambiguities or contradictions. Research by Dalpiaz et al. [32] demonstrates how semantic analysis can improve requirement clarity, helping teams identify missing or conflicting details early in the development cycle.
Automated Documentation.
Natural language generation can generate or update documentation based on source code changes. Moreno et al. [95] show that although such automation saves significant time, completeness and accuracy persist, underscoring the need for ongoing human oversight and validation.
Test Case Generation.
LLMs can convert requirement statements or code snippets into test scenarios, enhancing test coverage and catching edge cases. DeepTest [139] exemplifies how AI-based approaches can automatically craft test cases, notably for complex domains such as autonomous driving software.
Defect Detection and Code Review.
LLM-based models can flag potential security vulnerabilities, performance bottlenecks, or logical errors when integrated into code review pipelines. By drawing on learned patterns from large code repositories, these models offer a second layer of scrutiny, though explainability and false-positive rates remain concerns.
Compliance and Regulatory Checks.
Organizations bound by regulations (e.g., HIPAA, GDPR) can leverage LLMs to parse compliance documents and compare them against project artifacts. Discrepancies, such as missing user consent clauses or inadequate data handling procedures, can be highlighted for further investigation.
III-C Notable LLM Frameworks and Tools
OpenAI API Ecosystem.
OpenAI provides endpoints for code generation, text classification, and embedding-based similarity checks. Enterprise-ready features, including role-based access and usage analytics, cater to organizations seeking scalable AI solutions.
Microsoft’s Azure OpenAI Services.
Microsoft extends OpenAI’s models within Azure, offering compliance-centric features like data encryption at rest, secured endpoints, and traceable audit logs. These services facilitate LLM adoption in sectors with strict regulatory requirements.
Open-Source Tools.
Beyond commercial offerings, open-source efforts like CodeT5 [117], PolyCoder [158], and community-driven QA platforms integrate LLM-based bug detection, code completion, and textual analysis. They provide a customizable foundation for teams aiming to build or extend AI-driven quality assurance pipelines without relying on black-box vendor solutions.
IV Selection Criteria
We established structured selection criteria to ensure our survey remained focused and meaningful at the intersection of Large Language Models (LLMs) and Software Quality Assurance (SQA). We designed these criteria to align with industry-relevant quality goals, demonstrate technical depth, and capture the evolving role of LLMs in software engineering. We included a study in our review if it satisfied the following conditions:
  1. 1.
    Primary Focus on LLM Use in Software Quality Contexts: The study investigated the use of LLMs in tasks directly related to software quality assurance, such as requirement analysis, test generation, code quality improvement, bug detection, documentation enhancement, or standards compliance.
  2. 2.
    Alignment with SQA Objectives or Standards: While explicit references to formal standards (e.g., ISO/IEC 25010, ISO 12207, TMM, or CMMI) were preferred, studies were also included if they demonstrated alignment with core SQA objectives—such as reliability, maintainability, security, test coverage, or suitability —as defined by these frameworks.
  3. 3.
    Empirical, Technical, or Methodological Rigor: We prioritized papers that included empirical validation (e.g., case studies, benchmarks, experiments) or detailed methodology (e.g., tool architecture, model design, prompt engineering strategies) to ensure actionable insights and technical credibility.
  4. 4.
    Recency and Technological Relevance: Considering the rapid advancement of both LLMs and AI-in-SQA tools, we limited our survey to work published from 2023 onward, capturing the post-ChatGPT wave of innovation. We made exceptions for highly cited or foundational papers.
  5. 5.
    Breadth of SQA Activities Represented: To reflect the diversity of modern SQA, we included papers across multiple quality dimensions—functional correctness, test automation, security auditing, static analysis, configuration validation, and maintainability. This breadth allowed us to evaluate LLM applicability across the full SQA lifecycle.
  6. 6.
    Practical Relevance or Deployment Context: We preferred studies that demonstrated industrial relevance, provided tool availability, or offered deployment insights (e.g., GitHub Copilot evaluations, Meta’s testing infrastructure) to bridge academic insight with practice.
  7. 7.
    Peer-Reviewed or Preprint Research: Selected works were drawn from peer-reviewed journals, conference proceedings, or widely recognized preprint repositories (e.g., arXiv). This criterion ensured that all studies met a basic academic rigor and public dissemination standard.
V Characteristics of the Surveyed Literature
To characterize the state of the art in applying large language models (LLMs) to software quality assurance (SQA), we analyzed over 130 papers published between 2023 and early 2025. Below, we present key findings drawn from both quantitative distributions and qualitative inspection of representative studies.
V-A Publication Trends
As shown in Figure 1, the Number of papers surged dramatically in 2024, with over 85 publications—over double the output of 2023. This sharp growth underscores the burgeoning interest in LLMs for SQA and the availability of more capable open and commercial models during this period. The drop observed in 2025 is expected, given that the year is still in progress at the time of this writing.
Figure 1: Number of papers published per year from 2023 to early 2025, showing a rise in 2024 as interest in LLMs for software quality assurance surged.
V-B Dataset Utilization
Figure 2 reveals that nearly 30% of the reviewed papers did not specify any dataset or did not use one at all—highlighting a concerning gap in reproducibility. Among those that did, common dataset types included:
  • Open-source software projects, such as GitHub issues and repositories.
  • Benchmark suites like Defects4J and QuixBugs, particularly popular in testing applications.
  • Security and vulnerability datasets, as used in works like LLMs for Intelligent Software Testing: A Comparative Study.
Synthetic and proprietary datasets also appeared but were relatively rare. The use of domain-specific datasets (e.g., medical, legal) remains limited, indicating a focus on general-purpose or software-specific corpora.
Figure 2: Distribution of dataset themes used in the surveyed literature. A significant portion of papers did not specify any dataset, while open-source projects and benchmark suites were most common among those that did.
V-C Evaluation Approaches
Various evaluation methods were employed, as visualized in Figure 3. Common practices included:
  • Comparative evaluations, e.g., comparing LLM outputs against rule-based or traditional baselines.
  • Empirical/user studies, as seen in Advancing Requirements Engineering Through Generative AI, which collected practitioner feedback.
  • Automated performance metrics, such as precision, recall, BLEU, and accuracy.
A smaller but important subset used ablation analyses, static/dynamic analysis tools, or model-based frameworks.
Figure 3: Frequency of evaluation approaches used in the papers. Comparative studies, empirical/user evaluations, and automated performance metrics dominated the landscape.
V-D Fine-Tuning Adoption
As shown in Figure 4, only 14.3% of the studies reported using fine-tuned models. Most leveraged zero-shot or few-shot prompting on pre-trained models like ChatGPT or GPT-4. One exception is Requirements are All You Need: From Requirements to Code with LLMs, which fine-tunes on domain-specific data to improve generation accuracy.
Figure 4: Proportion of papers that used fine-tuned LLMs versus those that relied solely on pre-trained models. Most studies avoided fine-tuning.
V-E LLMs in Use
Figure 5 shows that GPT-4 was the most commonly used model (21%), followed by GPT-3.5, ChatGPT, and CodeT5. Open-source models like LLaMA and CodeLLaMA were used in academic or experimental studies. Nearly 19% of papers did not specify which LLM was used, raising reproducibility concerns.
Figure 5: Distribution of LLMs reported in the literature. GPT-4, GPT-3.5, and ChatGPT were the most commonly used, though many papers did not specify the model used.
V-F Prompting Strategies
Prompting strategies varied significantly (see Figure 6). Few-shot prompting was most popular (23.3%), followed by chain-of-thought (16.7%) and zero-shot (10%). A notable portion (14.4%) used “other” techniques, covering custom or hybrid strategies. More advanced techniques, such as instruction prompting, prompt chaining, and retrieval-augmented generation, were occasionally employed.
Figure 6: Distribution of prompting techniques employed across papers. Few-shot and chain-of-thought prompting were most prevalent, with some papers using custom or hybrid strategies.
In summary, the literature on LLMs in SQA is expanding rapidly. Most studies rely on general-purpose LLMs and prompt-based interaction rather than fine-tuning. Evaluation practices are becoming more diverse and rigorous, though dataset usage and reporting gaps remain.
VI Mapping LLM-based SQA Approaches to Established Standards
This section demonstrates how Large Language Model (LLM) based software quality assurance (SQA) methods can complement and extend existing frameworks and standards. Table I outlines representative applications of LLMs in the context of established SQA practices. By aligning AI-driven approaches with recognized industry guidelines, organizations can more effectively integrate LLMs into their quality assurance workflows while maintaining regulatory and methodological compliance.
TABLE I: Extended LLM-SQA Standards Mapping (Per Application)
Standard Quality Attribute/Process LLM Application Key References
ISO/IEC 12207 Software Requirements Translate stakeholder needs into requirement specifications [77, 152, 7, 91, 90, 154, 30]
Architecture and Design Recommend modular architecture and design patterns [154, 159, 71, 29, 7]
Implementation Generate code templates and enforce standards [46, 120, 113, 112, 155, 159, 71]
Verification Generate unit and integration tests [2, 9, 81, 125, 57, 15, 46, 102, 50]
Validation Compare test outcomes against requirements [50, 118, 15, 2, 9]
Configuration Management Track revisions and suggest updates [84, 83, 153, 149, 114, 129, 146]
ISO/IEC 25010 Functional Suitability Generate requirement-aligned test cases [77, 152, 9, 2, 91, 90, 30, 118, 15, 125, 102, 57]
Reliability Detect unstable patterns and improve fault tolerance [46, 50, 61, 107, 5, 116]
Usability Review and enhance UI/UX clarity and accessibility [43, 109, 41, 42, 19, 75, 89, 55, 164, 137]
Performance Efficiency Optimize inefficient code constructs [145, 26, 151, 52, 31, 122, 28, 99]
Maintainability Recommend refactorings and reduce code smells [159, 29, 10, 112, 165, 155, 113, 71, 101]
Portability Ensure cross-platform compatibility [149, 3, 84, 153, 83, 146]
Compatibility Analyze integration and dependency issues [120, 84, 88, 166, 114, 129]
Security Flag insecure practices and recommend secure alternatives [82, 168, 160, 161, 1, 63, 141, 124, 59, 78, 156, 17, 100, 106, 33]
ISO/IEC 5055 Reliability (Code-Level) Detect error-prone constructs, support fault localization [50, 61, 107, 9, 15, 46, 102]
Performance (Code-Level) Detect inefficiencies, optimize performance [145, 26, 52, 28, 31, 151, 99, 122, 144]
Security (Code-Level) Security audits, vulnerability detection [82, 168, 160, 161, 63, 1, 59, 78, 141, 106, 156, 100, 17, 33]
Maintainability (Code-Level) Detect low maintainability traits, suggest refactoring [159, 29, 10, 112, 165, 155, 113, 71, 101]
ISO 9001/9003 Customer Focus Extract insights from customer feedback [96, 162, 85, 150, 47, 131, 157]
Leadership Draft policies and quality objectives [7, 44, 48, 25]
People Engagement Generate onboarding and training content [108, 74]
Process Approach Standardize process documentation [167, 14, 76, 93]
Continuous Improvement Analyze historical data and recommend refinements [12, 13, 134]
Evidence-Based Decisions Summarize metrics and logs [56, 80, 23, 163, 70, 37, 135, 98]
Relationship Management Analyze communication logs [86, 22]
CMMI 2.0 Governance Support audits and risk assessments [7, 44, 48, 25, 8, 133, 58]
Operations Analyze logs and process inefficiencies [129, 153, 88, 120, 84, 149, 114, 3, 166]
Support Create SOPs and internal guides [108, 74, 97, 34, 14, 159, 155, 113]
Strategic Planning Summarize trends and support forecasting [132, 6, 56, 80, 163, 70, 22, 134]
TMM Level 1 - Initial Generate baseline unit tests [57, 81, 18, 126, 11]
Level 2 - Defined Create structured test plans and traceability [77, 152, 9, 90, 125]
Level 3 - Integrated Integrate tests with CI/CD pipelines [46, 102, 118, 30, 50]
Level 4 - Managed Analyze test metrics and coverage [2, 15, 126, 81, 9]
Level 5 - Optimized Recommend test optimization strategies [50, 46, 125, 11, 87, 110]
VI-A Aligning with ISO/IEC 12207
ISO/IEC 12207:2017 establishes a comprehensive framework for software life cycle processes, encompassing development, verification, validation, and maintenance [68]. LLM-based tools contribute across these stages by automating analysis, generation, and consistency checking throughout the software life cycle.
Software Requirements.
LLMs assist in translating unstructured stakeholder needs into structured requirement specifications, promoting clarity, traceability, and completeness. These tools can also detect ambiguities and inconsistencies during early-stage elicitation. For instance, Krishna et al. [77] demonstrated that GPT-4 could generate Software Requirements Specifications (SRS) comparable to entry-level engineers, effectively identifying and rectifying issues within requirements documents. Wei [152] explored the direct translation of requirements into code using LLMs, highlighting their potential to streamline the development process. Arora et al. [7] conducted a SWOT analysis, emphasizing LLMs’ strengths in requirements elicitation and validation.
Additionally, Lutze and Waldhör [91] showcased the use of LLMs in generating specifications from requirements documents for smart devices, enhancing the automation of the specification process. Lubos et al. [90] evaluated LLMs’ capabilities in assessing the quality of software requirements, aligning with ISO 29148 standards. White et al. [154] introduced prompt patterns to improve requirements elicitation using ChatGPT. Couder et al. [30] investigated requirements verification through source code analysis using LLMs.
Architecture and Design.
LLMs can recommend modular and maintainable architectural structures by analyzing functional decomposition and quality attributes. This includes proposing design patterns or modular breakdowns aligned with functional and non-functional requirements. White et al. [154] presented prompt patterns that guide LLMs in generating architectural designs and refactoring suggestions. Xu et al. [159] introduced MANTRA, a framework that enhances automated method-level refactoring through contextual retrieval-augmented generation and multi-agent LLM collaboration. Jelodar et al. [71] provided a comprehensive study on LLMs for source code analysis, discussing their applications in architectural design. Cordeiro et al. [29] conducted an empirical research on the code refactoring capabilities of LLMs, highlighting their effectiveness in improving software architecture. Arora et al. [7] emphasized LLMs’ role in advancing requirements engineering, including architectural considerations.
Implementation.
LLMs facilitate code synthesis from formal or informal specifications, enforce best practices, and perform AI-powered code reviews to ensure conformance to organizational standards. Fakhoury et al. [46] explored LLM-based test-driven interactive code generation, demonstrating their utility in producing code aligned with specifications. Rasheed et al. [120] discussed AI-powered code reviews with LLMs, highlighting early results in improving code quality. Pomian et al. [113] introduced EM-Assist, a safe automated extraction method refactoring tool using LLMs. Further, Pomian et al. [112] examined the integration of LLMs with IDE static analysis for extract method refactoring. Wu et al. [155] presented iSmell, which combines LLMs with expert toolsets for code smell detection and refactoring. Xu et al. [159] and Jelodar et al. [71] also contributed to understanding LLMs in code implementation and analysis.
Verification.
LLMs generate unit and integration tests from requirement documents and source code, accelerating early defect detection and increasing test coverage at lower cost. Alagarsamy et al. [2] enhanced LLMs for text-to-test case generation, improving test automation. Arora et al. [9] conducted an industrial study on generating test scenarios from natural language requirements using retrieval-augmented LLMs. Li et al. [81] evaluated LLMs for software testing, assessing their effectiveness in test generation. Sami et al. [125] developed a test case scenario generation tool using LLMs. Guilherme and Vincenzi [57] investigated ChatGPT’s capability in unit test generation. Bhatia et al. [15] discussed system test case design from requirements specifications using ChatGPT. Fakhoury et al. [46] and Pan et al. [102] further explored LLMs in test generation. Foster et al. [50] introduced mutation-guided LLM-based test generation at Meta.
Validation.
By comparing test execution logs against expected behaviors derived from requirements, LLMs help confirm that the system meets stakeholder intent. They also support post-test analysis and bug localization. Foster et al. [50] utilized mutation-guided LLM-based test generation to validate software behavior. Rahman and Zhu [118] automated user story generation with test case specification using LLMs. Bhatia et al. [15] highlighted using ChatGPT in system test case design for validation purposes. Alagarsamy et al. [2] and Arora et al. [9] also contributed to understanding LLMs in software validation.
Configuration Management.
LLMs support software configuration tasks such as tracking document revisions, identifying configuration drift, and suggesting updates across baselines, libraries, and API integrations. Lian et al. [84, 83] explored using LLMs as configuration validators, assessing their effectiveness in maintaining software configurations. Wen et al. [153] discussed LLM-based misconfiguration detection for AWS serverless computing. Wang et al. [149] identified performance-sensitive configurations in software systems through code analysis with LLM agents. Pornprasit and Tantithamthavorn [114] examined fine-tuning and prompt engineering for LLM-based code review automation. Shan et al. [129] introduced a two-stage strategy to localize configuration errors via logs using LLMs. Wang et al. [146] evaluated the facilitation of network configuration by LLMs.
VI-B Aligning with ISO/IEC 25010
The ISO/IEC 25010:2011 standard defines eight key product quality attributes: functionality, reliability, usability, performance efficiency, maintainability, portability, compatibility, and security [65]. LLMs (Large Language Models) have demonstrated capabilities that directly contribute to these attributes.
Functional Suitability
LLMs enable the automated derivation of test cases, acceptance criteria, and functional specifications from natural language requirements. Krishna et al. [77] empirically demonstrate LLMs’ ability to handle real-world requirements documents, while Wei [152] illustrates full pipeline support from requirements to code. As shown by Arora et al. [9], retrieval-augmented generation and tailored prompt engineering techniques enhance coverage and precision. Works by Alagarsamy et al. [2] and Lutze and Waldhör [91] further validate LLM effectiveness across industries, emphasizing gains in consistency and automation.
Reliability
LLMs support fault localization, code validation, and resilience by enabling test-driven development, as shown in empirical work by Fakhoury et al. [46]. Foster et al. [50] highlight improved mutation coverage in industrial settings at Meta. Hu et al. [61] caution against misuse, underlining the need for interpretability, while Pelliccione and Laranjeiro [107] contextualize LLM adoption trends in reliability engineering. Alshahwan et al. [5] propose strategies for offline robustness validation.
Usability
Duan et al. [43, 41] explore how LLMs assess and critique UI mockups using human-centered datasets. Petridis et al. [109] and Ghosh et al. [55] show LLMs generating feedback across domains, including healthcare interfaces. Zhang et al. [164] emphasize functional improvement in radiological systems, and Lu et al. [89] propose autonomous usability testing agents. These insights confirm LLMs can substantially enhance design coherence and accessibility.
Performance Efficiency
LLMs aid in detecting inefficient patterns and suggest optimized alternatives. Wadhwa et al. [145] show automated resolution of quality issues, while Wei et al. [151] present agent-based systems for optimizing parallelism. Comparative studies by Cui et al. [31] and Rosas et al. [122] evaluate LLM-generated code against classical compilers. Coignion et al. [28] assess runtime implications, and Gao et al. [52] advocate for search-based tuning of LLM-generated solutions.
Maintainability
LLMs support refactoring recommendations like the Extract Method and Move Method. Xu et al. [159] integrate contextual retrieval with multi-agent systems for automated method-level improvements. Cordeiro et al. [29], and Pomian et al. [112] demonstrate that LLMs outperform static analysis in maintainability-related tasks. Wu et al. [155] and Zhang et al. [165] show improvements in code smell reduction and method relocation. Jelodar et al. [71] consolidate source code analysis insights.
Portability
Wang et al. [149], and Albuquerque et al. [3] show LLMs can detect environment-specific configurations and suggest alternatives to enhance portability. Lian et al. [84] and Wen et al. [153] present tools for validating cross-platform readiness, including for AWS serverless setups. These studies demonstrate the ability of LLMs to generalize across environments empirically through benchmarked misconfiguration correction and support for compilation error tracing.
Compatibility
LLMs support static and dynamic validation of configuration consistency across environments. Lian et al. [84, 83] illustrate effective configuration validators. Pornprasit and Tantithamthavorn [114] show fine-tuned models outperform default LLMs for code review. Shan et al. [129] and Zhang et al. [166] demonstrate log-driven diagnostics and comment analysis, highlighting their compatibility assurance potential.
Security
LLMs detect and assist in resolving vulnerabilities through static and semantic analysis. Li et al. [82] and Zibaeirad and Vieira [168] explore LLMs in zero-shot and reinforcement learning-based detection scenarios. Yang et al. [160, 161] incorporate development context for fix suggestions. Studies by Islam et al. [63] and Guo et al. [59] emphasize LLMs’ adaptability across attack types. Tools like InferFix [72] and the SMART framework [17] validate LLM-based repair effectiveness. Pearce et al. [106] and de-Fitero-Dominguez et al. [33] reinforce the practical relevance of these techniques in software security pipelines.
VI-C Aligning with ISO/IEC 5055
ISO/IEC 5055:2021 defines a standardized framework for measuring structural code quality through four dimensions: Reliability, Performance Efficiency, Security, and Maintainability. LLM-based tools contribute to these through intelligent static analysis, code transformation, and vulnerability detection.
Reliability (Code-Level)
LLMs identify patterns correlated with reliability defects, such as unhandled exceptions or misuse of synchronization constructs. Foster et al. [50] show that mutation-guided LLM test generation increases fault detection in large industrial codebases. Fakhoury et al. [46] and Pan et al. [102] validate that LLMs generate unit tests covering boundary and edge cases. Hu et al. [61] warn about overreliance and stress proper tooling contexts, while Pelliccione and Laranjeiro [107] contextualize LLM reliability capabilities in safety-critical environments. Arora et al. [9] and Bhatia et al. [15] contribute practical strategies for integrating LLMs into fault-tolerant requirement validation pipelines.
Performance (Code-Level)
LLMs can profile and refactor performance bottlenecks. Wadhwa et al. [145, 144] highlight the resolution of code inefficiencies like nested loops or memory leaks. Wei et al. [151] demonstrate performance tuning of parallel applications using agent-driven optimization. Gao et al. [52] introduce a search-based framework to refine code using LLM feedback iteratively. Comparative benchmarks by Cui et al. [31], Niu et al. [99], and Rosas et al. [122] assess runtime performance and energy consumption, showing LLM-assisted improvements rival compiler-optimized code. Coignion et al. [28] quantify execution gains using LLM-generated LeetCode solutions.
Security (Code-Level)
LLMs support vulnerability detection by combining syntax, semantics, and contextual understanding. Li et al. [82] use LLMs for static vulnerability detection. Zibaeirad and Vieira [168] show zero-shot performance in identifying code weaknesses. Yang et al. [161, 160] integrate development history into vulnerability detection. Islam et al. [63] and Guo et al. [59] show adaptive reasoning across software stacks. Kulsum et al. [78] evaluate LLM patch accuracy and feedback loops. Pearce et al. [106], Noever [100], and De-Fitero-Dominguez et al. [33] assess LLM repair capabilities across benchmark datasets. These tools augment, not replace, static and dynamic analysis workflows.
Maintainability (Code-Level)
LLMs help refactor legacy or poorly structured code by identifying low maintainability features like code duplication, long methods, or high cyclomatic complexity. Xu et al. [159] integrate multi-agent collaboration for Extract Method refactoring. Cordeiro et al. [29], and Jelodar et al. [71] measure maintainability gains across open-source repositories. Pomian et al. [112, 113] propose hybrid models combining IDE tools and LLMs. Zhang et al. [165] introduce a deep learning framework for Move Method recommendations. Wu et al. [155] combines static tools with LLM insights to eliminate code smells, while Nunes et al. [101] assesses effectiveness in real-world systems. Electronics-based studies [10] show LLMs help identify and reduce structural anti-patterns such as data clumps.
VI-D Aligning with ISO 9001 and ISO/IEC 90003
ISO 9001:2015 establishes a framework for quality management across industries, while ISO/IEC 90003:2014 provides specific guidance for software engineering organizations. LLM-based systems can help organizations meet these standards through structured documentation, automated analysis, and process alignment.
Customer Focus
LLMs can extract structured insights from large volumes of customer feedback, enabling the identification of satisfaction drivers, pain points, and emerging themes. InsightNet [96] offers a structured pipeline for feedback analysis. Zhang et al. [162] explore LLMs as a conversational interface to engage with enterprise-scale feedback repositories. Falatouri et al. [47] and Soni [131] demonstrate improvements in customer lifecycle visibility, while Lin et al. [85] focus on satisfaction estimation using interpretable models. Wulf and Meierhofer [157] assess LLMs in automated service quality.
Leadership
Generative AI tools are increasingly used for strategic policy drafting and quality governance. Arora et al. [7] highlight their role in requirements alignment and documentation practices. Dzeparoska et al. [44] propose LLMs for intent-based application management, while Feng et al. [48] introduce collaborative policy prototyping frameworks. Cheong et al. [25] emphasize responsible AI integration in leadership-facing legal tools.
People Engagement
LLMs support disseminating organizational knowledge by generating tailored onboarding guides, technical wikis, and learning modules. Pereira et al. [108] evaluate their utility in education and training for software engineers. Kernan Freire et al. [74] benchmark LLM-powered knowledge-sharing systems in industrial manufacturing.
Process Approach
Consistent documentation of software processes is critical for quality assurance. Zhu et al. [167] present an LLM-driven approach to business process documentation. Kourani et al. [76] and Berti et al. [14] extend this to include process modeling and prompt-based process mining using LLMs. Mandvikar [93] illustrates enhanced workflows through intelligent document processing (IDP).
Continuous Improvement
LLMs analyze operational histories and retrospectives to support improvement cycles. Berti et al. [12, 13] demonstrate benchmarking strategies for evaluating LLMs in process mining contexts. Su et al. [134] propose a framework for documentation-driven feedback collection in retrospectives.
Evidence-Based Decisions
LLMs generate summaries and insights from structured and unstructured data, improving the speed and accuracy of decision-making. Tools like Godbole et al.’s long-context summarization framework [56] and Leiva-Araos et al.’s ensemble prompt method [80] exemplify scalable enterprise analytics. Other works focus on preprocessing [163], data pipelines [23], and anomaly detection for dashboards [37, 135, 70, 98].
Relationship Management
LLMs are used to mine sentiment, synthesize discussions, and surface emerging themes in stakeholder interactions. Liu and Sun [86] show their application in qualitative policy interviews, while Calderon and Reichart [22] trace interpretability trends relevant to stakeholder transparency.
Software Development Processes (ISO/IEC 90003)
LLMs accelerate the standardization of software practices through automated documentation synthesis and quality-driven use case alignment. Mandal et al. [92], and Tikayat et al. [140] explore synthesizing software specifications and agile engineering guidelines. Naimi et al. [97] and Della Porta et al. [34] document real-world applications in waterfall and hybrid life cycles. Birru [16] assesses their utility in agile technical documentation.
Maintenance (ISO/IEC 90003)
LLMs accelerate legacy code modernization through static analysis and refactoring support. Jelodar et al. [71] and Nunes et al. [101] evaluate maintainability improvements in industrial codebases. Diggs et al. [36] highlight practical modernization scenarios using LLM-generated documentation.
Supplier Evaluation (ISO/IEC 90003)
Vendor documentation and compliance artifacts are evaluated using LLM-powered tools. Arora et al. [8] and Garza et al. [54] show early frameworks for AI-assisted compliance checking. Doris et al. [39] and Singla et al. [130] examine vulnerability and documentation assessments in software supply chains. Sovrano et al. [133], Wang and Wu [148], Guldimann et al. [58], and Thummala and Rachaboyina [138] reinforce LLM utility in both technical and regulatory evaluation of external software dependencies.
VI-E Aligning with CMMI
Capability Maturity Model Integration (CMMI) provides a framework for enhancing software and organizational process maturity. LLMs can augment each phase of this model, from establishing basic governance practices to informing high-level strategic decisions.
Governance
LLMs contribute to governance by generating policies, supporting regulatory alignment, and automating compliance documentation. Arora et al. [7] demonstrate generative AI’s role in aligning requirements with governance objectives. Dzeparoska et al. [44] introduce LLM-based intent management for policy standardization. In contrast, Feng et al. [48] and Cheong et al. [25] explore collaborative policy development for legal and ethical compliance. In regulated domains, works like Sovrano et al. [133] and Guldimann et al. [58] evaluate LLM-driven compliance with emerging AI governance frameworks. Arora et al. [8] also emphasize LLMs in standards-conformant specification generation.
Operations
Operational support includes configuration validation, log-based diagnostics, and quality assurance. Shan et al. [129] propose log-driven strategies to localize configuration errors. Wen et al. [153] address misconfiguration detection in cloud-based environments, and Rasheed et al. [120] benchmark LLMs in code review and operational inspection tasks. Lu et al. [88] use fine-tuning techniques for operational code review, while Lian et al. [84] and Pornprasit et al. [114] explore runtime environment validation. Wang et al. [149], and Albuquerque et al. [3] confirm LLMs’ ability to detect runtime risks and configuration mismatches in complex systems. Zhang et al. [166] integrates LLMs with comment analysis to identify documentation mismatches in operational code.
Support
LLMs streamline knowledge transfer by generating technical documentation and training artifacts. Pereira et al. [108] evaluate their use in education and training for engineering roles. Kernan et al. [74] demonstrate LLM-driven documentation in industrial settings. Naimi et al. [97] and Della Porta et al. [34] present tools for process documentation and lifecycle coverage. Berti et al. [14] apply LLMs to process abstraction and modeling. Xu et al. [159] and Pomian et al. [113] showcase automated refactoring as part of internal toolchains, while Wu et al. [155] address maintainability support and documentation consistency.
Strategic Planning
For leadership and planning, LLMs extract trends, forecast risks, and summarize large datasets into actionable intelligence. Soru and Marshall [132] and Alzapiedi and Bihl [6] detail trend extraction using LLMs on temporal and semantic datasets. Godbole et al. [56] present multi-document summarization for enterprise intelligence, while Leiva-Araos et al. [80] evaluate ensemble prompting for scalable reporting. Jansen et al. [70] emphasize automation of data analysis pipelines, and Zhang et al. [163] demonstrate preprocessing and anomaly detection. Calderon and Reichart [22] link stakeholder sentiment to planning decisions, and Su et al. [134] explore how LLMs inform strategy via retrospective synthesis and knowledge surfacing.
VI-F Aligning with TMM
The Test Maturity Model (TMM) defines a roadmap for organizations to evolve their testing practices from ad hoc approaches to optimized, data-driven testing processes [21]. LLMs contribute meaningfully across all five maturity levels through test generation, coverage analysis, and intelligent automation.
Initial (Level 1)
At the earliest maturity level, where testing practices are largely unstructured, LLMs help establish a baseline by generating unit tests from natural language comments or code snippets. Guilherme and Vincenzi [57] evaluate ChatGPT’s ability to produce syntactically valid unit tests. Boukhlif et al. [18] offer comparative insights on LLM-generated test quality across models. Santos et al. [126] and Bayri and Demirel [11] demonstrate early adoption scenarios where LLMs seed the testing process in previously untested systems. Li et al. [81] benchmark LLM test generation across complexity levels, establishing confidence in foundational coverage.
Defined (Level 2)
As testing becomes more procedural, LLMs assist in formalizing documentation and establishing consistent test artifacts. Krishna et al. [77] and Wei [152] demonstrate LLMs aligning requirements to test documentation. Lubos et al. [90] integrate LLMs into test plan verification. Arora et al. [9] illustrate traceability across requirement-to-test chains using RAG (retrieval-augmented generation) techniques. Sami et al. [125] propose tools that automate scenario-based test creation with LLM prompts.
Integrated (Level 3)
Testing is continuous and part of CI/CD workflows at this level. LLMs help automate test generation as code changes, ensuring up-to-date regression coverage. Fakhoury et al. [46] integrate test generation into developer-in-the-loop systems. Pan et al. [102] show multilingual unit test generation across platforms. Rahman and Zhu [118] present pipelines for automated user story and test creation. Couder et al. [30] highlight bi-directional verification between code and requirements. Foster et al. [50] evaluate mutation-guided test generation at scale in CI environments.
Managed (Level 4)
Data becomes central to decision-making at this stage. LLMs analyze metrics such as coverage gaps, defect density, and test flakiness. Alagarsamy et al. [2] propose LLM frameworks that balance coverage goals with test redundancy. Bhatia et al. [15] identify challenges in ChatGPT-generated system-level test scenarios, offering feedback loops for improvement. Santos et al. [126] emphasize dashboard generation for testing KPIs. Li et al. [81] show how model-driven test metrics can augment traditional static coverage reports. Arora et al. [9] contribute insights into tracing test effectiveness against dynamically shifting requirements.
Optimized (Level 5)
At the highest maturity level, organizations optimize testing through prediction and adaptation. Foster et al. [50] describe risk-based prioritization in mutation-aware LLM pipelines. Fakhoury et al. [46] and Sami et al. [125] demonstrate adaptive testing systems that learn from historical outcomes. Bayri and Demirel [11] review LLMs in advanced software QA methodologies. Automated tools like those in Lops et al. [87] assess test suite quality and self-update strategies. Plein et al. [110] validate LLM feasibility in test case generation from unstructured bug reports, supporting automated defect-centric testing.
VII Challenges, Limitations, and Risks
VII-A Data Privacy and Security
Risk of Data Exposure
One of the most pressing concerns is the potential exposure of proprietary or sensitive code when using public APIs. Such unintentional leaks could violate confidentiality agreements or data protection mandates [4]. Sensitive information embedded in code, such as API keys or personal data, may be inadvertently processed by third-party services.
Mitigation Strategies
Organizations can opt for on-premises deployment or secure private cloud solutions to retain internal control over data flow. Additionally, fine-tuning models on anonymized or obfuscated datasets reduces the risk of revealing sensitive information while allowing domain-specific improvements. This trade-off between privacy and model performance highlights the need for robust data governance frameworks [121].
VII-B Model Bias and Ethical Considerations
Training Data Bias
LLMs trained on large, uncurated datasets frequently inherit biases reflected in their training corpora [94]. In software contexts, such biases might manifest as limited coverage of diverse user requirements, or preferential treatment of certain coding patterns prevalent in the training data.
Ethical and Legal Implications
Automated quality assurance decisions can overlook domain-specific ethical or legal requirements, such as patient confidentiality in healthcare applications [51]. Over-reliance on LLM outputs could lead to decisions that fail to account for contextual nuances or local regulations. Consequently, human oversight remains essential to prevent ethically or legally problematic outcomes.
VII-C Explainability and Transparency
Black Box Concerns
Transformers and other large neural architectures often lack inherent interpretability, making it challenging to justify the reasoning behind a particular suggestion [123]. In high-stakes domains like finance, aerospace, or healthcare, stakeholders may be unwilling to adopt AI-driven QA without clearer insights into how recommendations are generated.
Potential Solutions
Research into explainable AI (XAI) seeks to bridge this gap by proposing techniques such as attention visualization, summarized reasoning chains, or supplementary symbolic analysis [40]. Although these methods add transparency, they can increase computational overhead and complexity, highlighting a tension between model performance and interpretability.
VII-D Resource Requirements and Scalability
Computational Demands
Many state-of-the-art LLMs boast tens or hundreds of billions of parameters, necessitating substantial processing power and memory. Running these models at scale can incur high hardware, energy, and maintenance costs- a barrier for small to medium-sized enterprises with limited IT budgets [128].
Cost–Benefit Analysis
Organizations must weigh the potential gains in QA productivity and defect reduction against the expenses of acquiring and maintaining the requisite infrastructure. Some teams adopt a hybrid approach, using lighter, distilled models for routine tasks while reserving larger, more expensive models for complex or mission-critical analyses.
VII-E Governance and Auditing
Traceability and Logging
Robust governance frameworks require detailed logs of AI-driven decisions and developer overrides, ensuring accountability in scenarios where LLM-suggested changes lead to production issues [119]. These logs also facilitate post-deployment reviews and retrospective analyses, essential for regulated industries like finance or healthcare.
Standards Compliance
Many industry-specific regulations mandate auditable processes for software changes [64]. As organizations integrate LLMs into QA pipelines, they must demonstrate compliance with relevant ISO standards, legal statutes, and sector-specific best practices. This necessitates thorough documentation of how models are trained, validated, and continuously monitored.
VIII Future Directions
The convergence of Large Language Models (LLMs) and Software Quality Assurance (SQA) has already demonstrated significant potential. However, further innovations and research areas remain ripe for exploration. This section highlights emerging directions in model adaptability, deployment architectures, multimodal analysis, standards evolution, and cross-industry applications.
VIII-A Adaptive and Continual Learning
Incremental Fine-tuning
As codebases evolve, static models may become outdated, missing recent changes or new coding conventions. A promising direction is incremental fine-tuning, where LLMs continuously learn from recent commits, bug reports, and updated libraries [136]. Such adaptability can preserve or even improve accuracy over time, provided organizations balance retraining costs with tangible improvements in defect detection and code generation.
Active Learning Paradigms
Active learning leverages human intervention to refine the model. For instance, when an LLM detects ambiguous or high-risk outputs, it can query developers for clarification. These labeled examples feed back into the model, enhancing its understanding of domain-specific patterns [127]. This iterative feedback loop aligns well with agile development cycles, enabling more responsive and context-aware QA.
VIII-B Federated or On-Premises LLM Deployments
Privacy-Focused Approaches
Federated learning offers a method to aggregate insights from multiple sources without transferring raw data to a centralized server [73]. Only model parameters are shared in this model, not the underlying code or sensitive information. This design can help meet stringent regulatory requirements in domains like healthcare or finance.
Industry Examples
Privacy-conscious sectors like banking have begun piloting secure, private AI deployments [45]. By conducting model training and inference on-premises or within tightly controlled cloud environments, organizations maintain compliance with confidentiality and data protection regulations while still benefiting from advanced LLM-based QA.
VIII-C Advanced Multimodal and Integrated SQA
Beyond Text and Code
While current SQA practices often focus on source code and natural language requirements, real-world systems encompass multiple data modalities. Future approaches may integrate UML diagrams, real-time performance metrics, and user feedback logs, providing a more holistic view of software quality [60].
Potential Gains
Multimodal analysis can detect complex architectural flaws or emergent behaviors that text- or code-only methods might miss. For instance, a spike in user-reported issues alongside an unusual surge in memory usage could indicate a regression in system design. By correlating these signals, LLM-driven QA tools could propose deeper, more targeted improvements earlier in the development life cycle.
VIII-D Standard Updates to Incorporate AI/LLM Methodologies
Call for AI-Specific Guidelines
Existing SQA standards, such as ISO/IEC 25010 and ISO/IEC 5055, provide robust frameworks for measuring software quality but do not explicitly address AI-based tooling. As LLMs permeate more QA processes, there is a growing need for dedicated guidelines or amendments to outline best practices for AI-based testing, validation, and compliance [147].
Regulatory Initiatives
Governmental and international bodies are increasingly aware of the impact of AI on safety and ethics. For instance, the proposed EU AI Act would require transparent documentation and rigorous risk assessments for high-stakes AI systems [115]. Incorporating these mandates into software quality standards will be critical to ensuring both innovation and accountability.
VIII-E Cross-Industry and Cross-Domain Extensions
Knowledge Transfer
Techniques pioneered in one sector—such as automotive or aerospace—can be adapted for others, accelerating AI adoption in less digitized industries [103]. For example, diagnostic methods that detect sensor data anomalies in autonomous vehicles could inform QA practices in industrial robotics or medical devices.
Open Research Questions
Many open questions remain about handling domain-specific terminologies, compliance norms, and testing workflows across different fields. Further investigation is needed to develop generalized methods that retain enough adaptability to suit varied regulatory landscapes and stakeholder requirements, enabling LLM solutions to scale seamlessly across diverse domains.

### v3

@@
 V Characteristics of the Surveyed Literature
-To characterize the state of the art in applying large language models (LLMs) to software quality assurance (SQA), we analyzed over 130 papers published between 2023 and early 2025.
+To characterize the state of the art in applying large language models (LLMs) to software quality assurance (SQA), we analyzed over 223 papers published between 2023 and 2025.
 Below, we present key findings drawn from both quantitative distributions and qualitative inspection of representative studies.
 V-A Publication Trends
-As shown in Figure 1, the Number of papers surged dramatically in 2024, with over 85 publications—over double the output of 2023.
-This sharp growth underscores the burgeoning interest in LLMs for SQA and the availability of more capable open and commercial models during this period.
-The drop observed in 2025 is expected, given that the year is still in progress at the time of this writing.
-Figure 1: Number of papers published per year from 2023 to early 2025, showing a rise in 2024 as interest in LLMs for software quality assurance surged.
+As illustrated in Figure 1, the number of publications increased sharply after 2023.
+The field saw 31 papers in 2023, followed by a substantial rise to 91 papers in 2024—nearly a threefold increase.
+This surge reflects the rapidly growing interest in applying LLMs to SQA, likely driven by the release of more capable open-source and commercial models during this period.
+In 2025, the count is 127 papers, indicating continued momentum and suggesting that research activity in this area remains strong and accelerating.
+Figure 1: Number of papers published per year from 2023 to 2025, showing consisted growth in interest in LLMs for software quality assurance.
 V-B Dataset Utilization
@@
 Evaluation practices are becoming more diverse and rigorous, though dataset usage and reporting gaps remain.
-VI Mapping LLM-based SQA Approaches to Established Standards
-This section demonstrates how Large Language Model (LLM) based software quality assurance (SQA) methods can complement and extend existing frameworks and standards.
-Table I outlines representative applications of LLMs in the context of established SQA practices.
-By aligning AI-driven approaches with recognized industry guidelines, organizations can more effectively integrate LLMs into their quality assurance workflows while maintaining regulatory and methodological compliance.
 TABLE I: Extended LLM-SQA Standards Mapping (Per Application)
 Standard Quality Attribute/Process LLM Application Key References
-ISO/IEC 12207 Software Requirements Translate stakeholder needs into requirement specifications [77, 152, 7, 91, 90, 154, 30]
-Architecture and Design Recommend modular architecture and design patterns [154, 159, 71, 29, 7]
-Implementation Generate code templates and enforce standards [46, 120, 113, 112, 155, 159, 71]
-Verification Generate unit and integration tests [2, 9, 81, 125, 57, 15, 46, 102, 50]
-Validation Compare test outcomes against requirements [50, 118, 15, 2, 9]
-Configuration Management Track revisions and suggest updates [84, 83, 153, 149, 114, 129, 146]
-ISO/IEC 25010 Functional Suitability Generate requirement-aligned test cases [77, 152, 9, 2, 91, 90, 30, 118, 15, 125, 102, 57]
-Reliability Detect unstable patterns and improve fault tolerance [46, 50, 61, 107, 5, 116]
-Usability Review and enhance UI/UX clarity and accessibility [43, 109, 41, 42, 19, 75, 89, 55, 164, 137]
-Performance Efficiency Optimize inefficient code constructs [145, 26, 151, 52, 31, 122, 28, 99]
-Maintainability Recommend refactorings and reduce code smells [159, 29, 10, 112, 165, 155, 113, 71, 101]
-Portability Ensure cross-platform compatibility [149, 3, 84, 153, 83, 146]
-Compatibility Analyze integration and dependency issues [120, 84, 88, 166, 114, 129]
-Security Flag insecure practices and recommend secure alternatives [82, 168, 160, 161, 1, 63, 141, 124, 59, 78, 156, 17, 100, 106, 33]
-ISO/IEC 5055 Reliability (Code-Level) Detect error-prone constructs, support fault localization [50, 61, 107, 9, 15, 46, 102]
-Performance (Code-Level) Detect inefficiencies, optimize performance [145, 26, 52, 28, 31, 151, 99, 122, 144]
-Security (Code-Level) Security audits, vulnerability detection [82, 168, 160, 161, 63, 1, 59, 78, 141, 106, 156, 100, 17, 33]
-Maintainability (Code-Level) Detect low maintainability traits, suggest refactoring [159, 29, 10, 112, 165, 155, 113, 71, 101]
-ISO 9001/9003 Customer Focus Extract insights from customer feedback [96, 162, 85, 150, 47, 131, 157]
-Leadership Draft policies and quality objectives [7, 44, 48, 25]
+ISO/IEC 12207 Software Requirements Translate stakeholder needs into requirement specifications [120, 231, 16, 143, 142, 236, 50, 155, 183, 220, 77, 117, 218, 6, 88]
+Architecture and Design Recommend modular architecture and design patterns [236, 243, 108, 49, 16, 208, 159, 84, 258, 85, 52, 44, 221]
+Implementation Generate code templates and enforce standards [66, 181, 172, 171, 237, 243, 108, 48, 93, 95, 83, 166, 211, 235, 252]
+Verification Generate unit and integration tests [4, 18, 126, 191, 86, 28, 66, 157, 71, 35, 255, 7, 206, 197, 139, 227, 29, 41, 156, 115]
+Validation Compare test outcomes against requirements [71, 179, 28, 4, 18, 259, 92, 195]
+Configuration Management Track revisions and suggest updates [130, 129, 233, 228, 173, 196, 224, 131, 214, 40, 80, 15, 248, 234]
+ISO/IEC 25010 Functional Suitability Generate requirement-aligned test cases [120, 231, 18, 4, 143, 142, 50, 179, 28, 191, 157, 86, 54]
+Reliability Detect unstable patterns and improve fault tolerance [66, 71, 91, 165, 11, 176, 14, 110, 190, 109, 241]
+Usability Review and enhance UI/UX clarity and accessibility [63, 168, 61, 62, 32, 116, 141, 78, 254, 207, 22, 68, 250, 161, 2, 125, 97]
+Performance Efficiency Optimize inefficient code constructs [223, 45, 230, 75, 51, 185, 47, 152, 247, 178, 105, 9]
+Maintainability Recommend refactorings and reduce code smells [243, 49, 23, 171, 256, 237, 172, 108, 154, 136, 132, 19, 160, 213, 244, 135, 10]
+Portability Ensure cross-platform compatibility [228, 5, 130, 233, 129, 224, 149]
+Compatibility Analyze integration and dependency issues [181, 130, 140, 257, 173, 196, 144, 79, 21]
+Security Flag insecure practices and recommend secure alternatives [128, 261, 245, 246, 3, 99, 215, 188, 89, 121, 238, 30, 153, 164, 55, 189, 14, 42]
+ISO/IEC 5055 Reliability (Code-Level) Detect error-prone constructs, support fault localization [71, 91, 165, 18, 28, 66, 157, 175, 242, 241, 201]
+Performance (Code-Level) Detect inefficiencies, optimize performance [223, 45, 75, 47, 51, 230, 152, 185, 222, 96, 9, 247]
+Security (Code-Level) Security audits, vulnerability detection [128, 261, 245, 246, 99, 3, 89, 121, 215, 164, 238, 153, 30, 55, 114, 209, 198, 212, 137]
+Maintainability (Code-Level) Detect low maintainability traits, suggest refactoring [243, 49, 23, 171, 256, 237, 172, 108, 154, 82, 184, 12, 37]
+ISO 9001/9003 Customer Focus Extract insights from customer feedback [148, 251, 133, 229, 67, 199, 239, 1, 232, 20, 249, 186, 122]
+Leadership Draft policies and quality objectives [260, 27, 118, 145, 72, 119]
 People Engagement Generate onboarding and training content [167, 113]
-Process Approach Standardize process documentation [167, 14, 76, 93]
+Process Approach Standardize process documentation [260, 27, 118, 145, 72, 119]
 Continuous Improvement Analyze historical data and recommend refinements [25, 26, 203]
 Evidence-Based Decisions Summarize metrics and logs [81, 124, 38, 253, 107, 58, 204, 151]
-Relationship Management Analyze communication logs [86, 22]
-CMMI 2.0 Governance Support audits and risk assessments [7, 44, 48, 25, 8, 133, 58]
-Operations Analyze logs and process inefficiencies [129, 153, 88, 120, 84, 149, 114, 3, 166]
-Support Create SOPs and internal guides [108, 74, 97, 34, 14, 159, 155, 113]
+Relationship Management Analyze communication logs [134, 36, 94]
+CMMI 2.0 Governance Support audits and risk assessments [16, 64, 69, 43, 17, 202, 87, 127]
+Operations Analyze logs and process inefficiencies [196, 233, 140, 181, 130, 228, 173, 5, 257, 73]
+Support Create SOPs and internal guides [167, 113, 150, 56, 27, 243, 237, 172, 226, 112]
 Strategic Planning Summarize trends and support forecasting [200, 13, 81, 124, 253, 107, 36, 203]
 TMM Level 1 - Initial Generate baseline unit tests [86, 126, 31, 192, 24]
-Level 2 - Defined Create structured test plans and traceability [77, 152, 9, 90, 125]
+Level 2 - Defined Create structured test plans and traceability [120, 231, 18, 142, 191, 219]
 Level 3 - Integrated Integrate tests with CI/CD pipelines [66, 157, 179, 50, 71]
@@
 Level 5 - Optimized Recommend test optimization strategies [71, 66, 191, 24, 138, 169]
-VI-A Aligning with ISO/IEC 12207
-ISO/IEC 12207:2017 establishes a comprehensive framework for software life cycle processes, encompassing development, verification, validation, and maintenance [68].
-LLM-based tools contribute across these stages by automating analysis, generation, and consistency checking throughout the software life cycle.
-Software Requirements.
-LLMs assist in translating unstructured stakeholder needs into structured requirement specifications, promoting clarity, traceability, and completeness.
-These tools can also detect ambiguities and inconsistencies during early-stage elicitation.
-For instance, Krishna et al.
-[77] demonstrated that GPT-4 could generate Software Requirements Specifications (SRS) comparable to entry-level engineers, effectively identifying and rectifying issues within requirements documents.
-Wei [152] explored the direct translation of requirements into code using LLMs, highlighting their potential to streamline the development process.
-Arora et al.
-[7] conducted a SWOT analysis, emphasizing LLMs’ strengths in requirements elicitation and validation.
-Additionally, Lutze and Waldhör [91] showcased the use of LLMs in generating specifications from requirements documents for smart devices, enhancing the automation of the specification process.
-Lubos et al.
-[90] evaluated LLMs’ capabilities in assessing the quality of software requirements, aligning with ISO 29148 standards.
-White et al.
-[154] introduced prompt patterns to improve requirements elicitation using ChatGPT.
-Couder et al.
-[30] investigated requirements verification through source code analysis using LLMs.
-Architecture and Design.
-LLMs can recommend modular and maintainable architectural structures by analyzing functional decomposition and quality attributes.
-This includes proposing design patterns or modular breakdowns aligned with functional and non-functional requirements.
-White et al.
-[154] presented prompt patterns that guide LLMs in generating architectural designs and refactoring suggestions.
-Xu et al.
-[159] introduced MANTRA, a framework that enhances automated method-level refactoring through contextual retrieval-augmented generation and multi-agent LLM collaboration.
-Jelodar et al.
-[71] provided a comprehensive study on LLMs for source code analysis, discussing their applications in architectural design.
-Cordeiro et al.
-[29] conducted an empirical research on the code refactoring capabilities of LLMs, highlighting their effectiveness in improving software architecture.
-Arora et al.
-[7] emphasized LLMs’ role in advancing requirements engineering, including architectural considerations.
-Implementation.
-LLMs facilitate code synthesis from formal or informal specifications, enforce best practices, and perform AI-powered code reviews to ensure conformance to organizational standards.
-Fakhoury et al.
-[46] explored LLM-based test-driven interactive code generation, demonstrating their utility in producing code aligned with specifications.
-Rasheed et al.
-[120] discussed AI-powered code reviews with LLMs, highlighting early results in improving code quality.
-Pomian et al.
-[113] introduced EM-Assist, a safe automated extraction method refactoring tool using LLMs.
-Further, Pomian et al.
-[112] examined the integration of LLMs with IDE static analysis for extract method refactoring.
-Wu et al.
-[155] presented iSmell, which combines LLMs with expert toolsets for code smell detection and refactoring.
-Xu et al.
-[159] and Jelodar et al.
-[71] also contributed to understanding LLMs in code implementation and analysis.
-Verification.
-LLMs generate unit and integration tests from requirement documents and source code, accelerating early defect detection and increasing test coverage at lower cost.
-Alagarsamy et al.
-[2] enhanced LLMs for text-to-test case generation, improving test automation.
-Arora et al.
-[9] conducted an industrial study on generating test scenarios from natural language requirements using retrieval-augmented LLMs.
-Li et al.
-[81] evaluated LLMs for software testing, assessing their effectiveness in test generation.
-Sami et al.
-[125] developed a test case scenario generation tool using LLMs.
-Guilherme and Vincenzi [57] investigated ChatGPT’s capability in unit test generation.
-Bhatia et al.
-[15] discussed system test case design from requirements specifications using ChatGPT.
-Fakhoury et al.
-[46] and Pan et al.
-[102] further explored LLMs in test generation.
-Foster et al.
-[50] introduced mutation-guided LLM-based test generation at Meta.
-Validation.
-By comparing test execution logs against expected behaviors derived from requirements, LLMs help confirm that the system meets stakeholder intent.
-They also support post-test analysis and bug localization.
-Foster et al.
-[50] utilized mutation-guided LLM-based test generation to validate software behavior.
-Rahman and Zhu [118] automated user story generation with test case specification using LLMs.
-Bhatia et al.
-[15] highlighted using ChatGPT in system test case design for validation purposes.
-Alagarsamy et al.
-[2] and Arora et al.
-[9] also contributed to understanding LLMs in software validation.
-Configuration Management.
-LLMs support software configuration tasks such as tracking document revisions, identifying configuration drift, and suggesting updates across baselines, libraries, and API integrations.
-Lian et al.
-[84, 83] explored using LLMs as configuration validators, assessing their effectiveness in maintaining software configurations.
-Wen et al.
-[153] discussed LLM-based misconfiguration detection for AWS serverless computing.
-Wang et al.
-[149] identified performance-sensitive configurations in software systems through code analysis with LLM agents.
-Pornprasit and Tantithamthavorn [114] examined fine-tuning and prompt engineering for LLM-based code review automation.
-Shan et al.
-[129] introduced a two-stage strategy to localize configuration errors via logs using LLMs.
-Wang et al.
-[146] evaluated the facilitation of network configuration by LLMs.
-VI-B Aligning with ISO/IEC 25010
-The ISO/IEC 25010:2011 standard defines eight key product quality attributes: functionality, reliability, usability, performance efficiency, maintainability, portability, compatibility, and security [65].
-LLMs (Large Language Models) have demonstrated capabilities that directly contribute to these attributes.
-Functional Suitability
-LLMs enable the automated derivation of test cases, acceptance criteria, and functional specifications from natural language requirements.
-Krishna et al.
-[77] empirically demonstrate LLMs’ ability to handle real-world requirements documents, while Wei [152] illustrates full pipeline support from requirements to code.
-As shown by Arora et al.
-[9], retrieval-augmented generation and tailored prompt engineering techniques enhance coverage and precision.
-Works by Alagarsamy et al.
-[2] and Lutze and Waldhör [91] further validate LLM effectiveness across industries, emphasizing gains in consistency and automation.
-Reliability
-LLMs support fault localization, code validation, and resilience by enabling test-driven development, as shown in empirical work by Fakhoury et al.
-[46].
-Foster et al.
-[50] highlight improved mutation coverage in industrial settings at Meta.
-Hu et al.
-[61] caution against misuse, underlining the need for interpretability, while Pelliccione and Laranjeiro [107] contextualize LLM adoption trends in reliability engineering.
-Alshahwan et al.
-[5] propose strategies for offline robustness validation.
-Usability
-Duan et al.
-[43, 41] explore how LLMs assess and critique UI mockups using human-centered datasets.
-Petridis et al.
-[109] and Ghosh et al.
-[55] show LLMs generating feedback across domains, including healthcare interfaces.
-Zhang et al.
-[164] emphasize functional improvement in radiological systems, and Lu et al.
-[89] propose autonomous usability testing agents.
-These insights confirm LLMs can substantially enhance design coherence and accessibility.
-Performance Efficiency
-LLMs aid in detecting inefficient patterns and suggest optimized alternatives.
-Wadhwa et al.
-[145] show automated resolution of quality issues, while Wei et al.
-[151] present agent-based systems for optimizing parallelism.
-Comparative studies by Cui et al.
-[31] and Rosas et al.
-[122] evaluate LLM-generated code against classical compilers.
-Coignion et al.
-[28] assess runtime implications, and Gao et al.
-[52] advocate for search-based tuning of LLM-generated solutions.
-Maintainability
-LLMs support refactoring recommendations like the Extract Method and Move Method.
-Xu et al.
-[159] integrate contextual retrieval with multi-agent systems for automated method-level improvements.
-Cordeiro et al.
-[29], and Pomian et al.
-[112] demonstrate that LLMs outperform static analysis in maintainability-related tasks.
-Wu et al.
-[155] and Zhang et al.
-[165] show improvements in code smell reduction and method relocation.
-Jelodar et al.
-[71] consolidate source code analysis insights.
-Portability
-Wang et al.
-[149], and Albuquerque et al.
-[3] show LLMs can detect environment-specific configurations and suggest alternatives to enhance portability.
-Lian et al.
-[84] and Wen et al.
-[153] present tools for validating cross-platform readiness, including for AWS serverless setups.
-These studies demonstrate the ability of LLMs to generalize across environments empirically through benchmarked misconfiguration correction and support for compilation error tracing.
-Compatibility
-LLMs support static and dynamic validation of configuration consistency across environments.
-Lian et al.
-[84, 83] illustrate effective configuration validators.
-Pornprasit and Tantithamthavorn [114] show fine-tuned models outperform default LLMs for code review.
-Shan et al.
-[129] and Zhang et al.
-[166] demonstrate log-driven diagnostics and comment analysis, highlighting their compatibility assurance potential.
-Security
-LLMs detect and assist in resolving vulnerabilities through static and semantic analysis.
-Li et al.
-[82] and Zibaeirad and Vieira [168] explore LLMs in zero-shot and reinforcement learning-based detection scenarios.
-Yang et al.
-[160, 161] incorporate development context for fix suggestions.
-Studies by Islam et al.
-[63] and Guo et al.
-[59] emphasize LLMs’ adaptability across attack types.
-Tools like InferFix [72] and the SMART framework [17] validate LLM-based repair effectiveness.
-Pearce et al.
-[106] and de-Fitero-Dominguez et al.
-[33] reinforce the practical relevance of these techniques in software security pipelines.
-VI-C Aligning with ISO/IEC 5055
-ISO/IEC 5055:2021 defines a standardized framework for measuring structural code quality through four dimensions: Reliability, Performance Efficiency, Security, and Maintainability.
-LLM-based tools contribute to these through intelligent static analysis, code transformation, and vulnerability detection.
-Reliability (Code-Level)
-LLMs identify patterns correlated with reliability defects, such as unhandled exceptions or misuse of synchronization constructs.
-Foster et al.
-[50] show that mutation-guided LLM test generation increases fault detection in large industrial codebases.
-Fakhoury et al.
-[46] and Pan et al.
-[102] validate that LLMs generate unit tests covering boundary and edge cases.
-Hu et al.
-[61] warn about overreliance and stress proper tooling contexts, while Pelliccione and Laranjeiro [107] contextualize LLM reliability capabilities in safety-critical environments.
-Arora et al.
-[9] and Bhatia et al.
-[15] contribute practical strategies for integrating LLMs into fault-tolerant requirement validation pipelines.
-Performance (Code-Level)
-LLMs can profile and refactor performance bottlenecks.
-Wadhwa et al.
-[145, 144] highlight the resolution of code inefficiencies like nested loops or memory leaks.
-Wei et al.
-[151] demonstrate performance tuning of parallel applications using agent-driven optimization.
-Gao et al.
-[52] introduce a search-based framework to refine code using LLM feedback iteratively.
-Comparative benchmarks by Cui et al.
-[31], Niu et al.
-[99], and Rosas et al.
-[122] assess runtime performance and energy consumption, showing LLM-assisted improvements rival compiler-optimized code.
-Coignion et al.
-[28] quantify execution gains using LLM-generated LeetCode solutions.
-Security (Code-Level)
-LLMs support vulnerability detection by combining syntax, semantics, and contextual understanding.
-Li et al.
-[82] use LLMs for static vulnerability detection.
-Zibaeirad and Vieira [168] show zero-shot performance in identifying code weaknesses.
-Yang et al.
-[161, 160] integrate development history into vulnerability detection.
-Islam et al.
-[63] and Guo et al.
-[59] show adaptive reasoning across software stacks.
-Kulsum et al.
-[78] evaluate LLM patch accuracy and feedback loops.
-Pearce et al.
-[106], Noever [100], and De-Fitero-Dominguez et al.
-[33] assess LLM repair capabilities across benchmark datasets.
-These tools augment, not replace, static and dynamic analysis workflows.
-Maintainability (Code-Level)
-LLMs help refactor legacy or poorly structured code by identifying low maintainability features like code duplication, long methods, or high cyclomatic complexity.
-Xu et al.
-[159] integrate multi-agent collaboration for Extract Method refactoring.
-Cordeiro et al.
-[29], and Jelodar et al.
-[71] measure maintainability gains across open-source repositories.
-Pomian et al.
-[112, 113] propose hybrid models combining IDE tools and LLMs.
-Zhang et al.
-[165] introduce a deep learning framework for Move Method recommendations.
-Wu et al.
-[155] combines static tools with LLM insights to eliminate code smells, while Nunes et al.
-[101] assesses effectiveness in real-world systems.
-Electronics-based studies [10] show LLMs help identify and reduce structural anti-patterns such as data clumps.
-VI-D Aligning with ISO 9001 and ISO/IEC 90003
-ISO 9001:2015 establishes a framework for quality management across industries, while ISO/IEC 90003:2014 provides specific guidance for software engineering organizations.
-LLM-based systems can help organizations meet these standards through structured documentation, automated analysis, and process alignment.
-Customer Focus
-LLMs can extract structured insights from large volumes of customer feedback, enabling the identification of satisfaction drivers, pain points, and emerging themes.
-InsightNet [96] offers a structured pipeline for feedback analysis.
-Zhang et al.
-[162] explore LLMs as a conversational interface to engage with enterprise-scale feedback repositories.
-Falatouri et al.
-[47] and Soni [131] demonstrate improvements in customer lifecycle visibility, while Lin et al.
-[85] focus on satisfaction estimation using interpretable models.
-Wulf and Meierhofer [157] assess LLMs in automated service quality.
-Leadership
-Generative AI tools are increasingly used for strategic policy drafting and quality governance.
-Arora et al.
-[7] highlight their role in requirements alignment and documentation practices.
-Dzeparoska et al.
-[44] propose LLMs for intent-based application management, while Feng et al.
-[48] introduce collaborative policy prototyping frameworks.
-Cheong et al.
-[25] emphasize responsible AI integration in leadership-facing legal tools.
-People Engagement
-LLMs support disseminating organizational knowledge by generating tailored onboarding guides, technical wikis, and learning modules.
-Pereira et al.
-[108] evaluate their utility in education and training for software engineers.
-Kernan Freire et al.
-[74] benchmark LLM-powered knowledge-sharing systems in industrial manufacturing.
-Process Approach
-Consistent documentation of software processes is critical for quality assurance.
-Zhu et al.
-[167] present an LLM-driven approach to business process documentation.
-Kourani et al.
-[76] and Berti et al.
-[14] extend this to include process modeling and prompt-based process mining using LLMs.
-Mandvikar [93] illustrates enhanced workflows through intelligent document processing (IDP).
-Continuous Improvement
-LLMs analyze operational histories and retrospectives to support improvement cycles.
-Berti et al.
-[12, 13] demonstrate benchmarking strategies for evaluating LLMs in process mining contexts.
-Su et al.
-[134] propose a framework for documentation-driven feedback collection in retrospectives.
-Evidence-Based Decisions
-LLMs generate summaries and insights from structured and unstructured data, improving the speed and accuracy of decision-making.
-Tools like Godbole et al.’s long-context summarization framework [56] and Leiva-Araos et al.’s ensemble prompt method [80] exemplify scalable enterprise analytics.
-Other works focus on preprocessing [163], data pipelines [23], and anomaly detection for dashboards [37, 135, 70, 98].
-Relationship Management
-LLMs are used to mine sentiment, synthesize discussions, and surface emerging themes in stakeholder interactions.
-Liu and Sun [86] show their application in qualitative policy interviews, while Calderon and Reichart [22] trace interpretability trends relevant to stakeholder transparency.
-Software Development Processes (ISO/IEC 90003)
-LLMs accelerate the standardization of software practices through automated documentation synthesis and quality-driven use case alignment.
-Mandal et al.
-[92], and Tikayat et al.
-[140] explore synthesizing software specifications and agile engineering guidelines.
-Naimi et al.
-[97] and Della Porta et al.
-[34] document real-world applications in waterfall and hybrid life cycles.
-Birru [16] assesses their utility in agile technical documentation.
-Maintenance (ISO/IEC 90003)
-LLMs accelerate legacy code modernization through static analysis and refactoring support.
-Jelodar et al.
-[71] and Nunes et al.
-[101] evaluate maintainability improvements in industrial codebases.
-Diggs et al.
-[36] highlight practical modernization scenarios using LLM-generated documentation.
-Supplier Evaluation (ISO/IEC 90003)
-Vendor documentation and compliance artifacts are evaluated using LLM-powered tools.
-Arora et al.
-[8] and Garza et al.
-[54] show early frameworks for AI-assisted compliance checking.
-Doris et al.
-[39] and Singla et al.
-[130] examine vulnerability and documentation assessments in software supply chains.
-Sovrano et al.
-[133], Wang and Wu [148], Guldimann et al.
-[58], and Thummala and Rachaboyina [138] reinforce LLM utility in both technical and regulatory evaluation of external software dependencies.
-VI-E Aligning with CMMI
-Capability Maturity Model Integration (CMMI) provides a framework for enhancing software and organizational process maturity.
-LLMs can augment each phase of this model, from establishing basic governance practices to informing high-level strategic decisions.
-Governance
-LLMs contribute to governance by generating policies, supporting regulatory alignment, and automating compliance documentation.
-Arora et al.
-[7] demonstrate generative AI’s role in aligning requirements with governance objectives.
-Dzeparoska et al.
-[44] introduce LLM-based intent management for policy standardization.
-In contrast, Feng et al.
-[48] and Cheong et al.
-[25] explore collaborative policy development for legal and ethical compliance.
-In regulated domains, works like Sovrano et al.
-[133] and Guldimann et al.
-[58] evaluate LLM-driven compliance with emerging AI governance frameworks.
-Arora et al.
-[8] also emphasize LLMs in standards-conformant specification generation.
-Operations
-Operational support includes configuration validation, log-based diagnostics, and quality assurance.
-Shan et al.
-[129] propose log-driven strategies to localize configuration errors.
-Wen et al.
-[153] address misconfiguration detection in cloud-based environments, and Rasheed et al.
-[120] benchmark LLMs in code review and operational inspection tasks.
-Lu et al.
-[88] use fine-tuning techniques for operational code review, while Lian et al.
-[84] and Pornprasit et al.
-[114] explore runtime environment validation.
-Wang et al.
-[149], and Albuquerque et al.
-[3] confirm LLMs’ ability to detect runtime risks and configuration mismatches in complex systems.
-Zhang et al.
-[166] integrates LLMs with comment analysis to identify documentation mismatches in operational code.
-Support
-LLMs streamline knowledge transfer by generating technical documentation and training artifacts.
-Pereira et al.
-[108] evaluate their use in education and training for engineering roles.
-Kernan et al.
-[74] demonstrate LLM-driven documentation in industrial settings.
-Naimi et al.
-[97] and Della Porta et al.
-[34] present tools for process documentation and lifecycle coverage.
-Berti et al.
-[14] apply LLMs to process abstraction and modeling.
-Xu et al.
-[159] and Pomian et al.
-[113] showcase automated refactoring as part of internal toolchains, while Wu et al.
-[155] address maintainability support and documentation consistency.
-Strategic Planning
-For leadership and planning, LLMs extract trends, forecast risks, and summarize large datasets into actionable intelligence.
-Soru and Marshall [132] and Alzapiedi and Bihl [6] detail trend extraction using LLMs on temporal and semantic datasets.
-Godbole et al.
-[56] present multi-document summarization for enterprise intelligence, while Leiva-Araos et al.
-[80] evaluate ensemble prompting for scalable reporting.
-Jansen et al.
-[70] emphasize automation of data analysis pipelines, and Zhang et al.
-[163] demonstrate preprocessing and anomaly detection.
-Calderon and Reichart [22] link stakeholder sentiment to planning decisions, and Su et al.
-[134] explore how LLMs inform strategy via retrospective synthesis and knowledge surfacing.
-VI-F Aligning with TMM
-The Test Maturity Model (TMM) defines a roadmap for organizations to evolve their testing practices from ad hoc approaches to optimized, data-driven testing processes [21].
-LLMs contribute meaningfully across all five maturity levels through test generation, coverage analysis, and intelligent automation.
-Initial (Level 1)
-At the earliest maturity level, where testing practices are largely unstructured, LLMs help establish a baseline by generating unit tests from natural language comments or code snippets.
-Guilherme and Vincenzi [57] evaluate ChatGPT’s ability to produce syntactically valid unit tests.
-Boukhlif et al.
-[18] offer comparative insights on LLM-generated test quality across models.
-Santos et al.
-[126] and Bayri and Demirel [11] demonstrate early adoption scenarios where LLMs seed the testing process in previously untested systems.
-Li et al.
-[81] benchmark LLM test generation across complexity levels, establishing confidence in foundational coverage.
-Defined (Level 2)
-As testing becomes more procedural, LLMs assist in formalizing documentation and establishing consistent test artifacts.
-Krishna et al.
-[77] and Wei [152] demonstrate LLMs aligning requirements to test documentation.
-Lubos et al.
-[90] integrate LLMs into test plan verification.
-Arora et al.
-[9] illustrate traceability across requirement-to-test chains using RAG (retrieval-augmented generation) techniques.
-Sami et al.
-[125] propose tools that automate scenario-based test creation with LLM prompts.
-Integrated (Level 3)
-Testing is continuous and part of CI/CD workflows at this level.
-LLMs help automate test generation as code changes, ensuring up-to-date regression coverage.
-Fakhoury et al.
-[46] integrate test generation into developer-in-the-loop systems.
-Pan et al.
-[102] show multilingual unit test generation across platforms.
-Rahman and Zhu [118] present pipelines for automated user story and test creation.
-Couder et al.
-[30] highlight bi-directional verification between code and requirements.
-Foster et al.
-[50] evaluate mutation-guided test generation at scale in CI environments.
-Managed (Level 4)
-Data becomes central to decision-making at this stage.
-LLMs analyze metrics such as coverage gaps, defect density, and test flakiness.
-Alagarsamy et al.
-[2] propose LLM frameworks that balance coverage goals with test redundancy.
-Bhatia et al.
-[15] identify challenges in ChatGPT-generated system-level test scenarios, offering feedback loops for improvement.
-Santos et al.
-[126] emphasize dashboard generation for testing KPIs.
-Li et al.
-[81] show how model-driven test metrics can augment traditional static coverage reports.
-Arora et al.
-[9] contribute insights into tracing test effectiveness against dynamically shifting requirements.
-Optimized (Level 5)
-At the highest maturity level, organizations optimize testing through prediction and adaptation.
-Foster et al.
-[50] describe risk-based prioritization in mutation-aware LLM pipelines.
-Fakhoury et al.
-[46] and Sami et al.
-[125] demonstrate adaptive testing systems that learn from historical outcomes.
-Bayri and Demirel [11] review LLMs in advanced software QA methodologies.
-Automated tools like those in Lops et al.
-[87] assess test suite quality and self-update strategies.
-Plein et al.
-[110] validate LLM feasibility in test case generation from unstructured bug reports, supporting automated defect-centric testing.
-VII Challenges, Limitations, and Risks
-VII-A Data Privacy and Security
+VI Challenges, Limitations, and Risks
+VI-A Data Privacy and Security
 Risk of Data Exposure
@@
 This trade-off between privacy and model performance highlights the need for robust data governance frameworks [182].
-VII-B Model Bias and Ethical Considerations
+VI-B Model Bias and Ethical Considerations
 Training Data Bias
@@
 Consequently, human oversight remains essential to prevent ethically or legally problematic outcomes.
-VII-C Explainability and Transparency
+VI-C Explainability and Transparency
 Black Box Concerns
@@
 Although these methods add transparency, they can increase computational overhead and complexity, highlighting a tension between model performance and interpretability.
-VII-D Resource Requirements and Scalability
+VI-D Resource Requirements and Scalability
 Computational Demands
@@
 Some teams adopt a hybrid approach, using lighter, distilled models for routine tasks while reserving larger, more expensive models for complex or mission-critical analyses.
-VII-E Governance and Auditing
+VI-E Governance and Auditing
 Traceability and Logging
@@
 This necessitates thorough documentation of how models are trained, validated, and continuously monitored.
-VIII Future Directions
+VII Future Directions
 The convergence of Large Language Models (LLMs) and Software Quality Assurance (SQA) has already demonstrated significant potential.
@@
 This section highlights emerging directions in model adaptability, deployment architectures, multimodal analysis, standards evolution, and cross-industry applications.
-VIII-A Adaptive and Continual Learning
+VII-A Adaptive and Continual Learning
 Incremental Fine-tuning
@@
 This iterative feedback loop aligns well with agile development cycles, enabling more responsive and context-aware QA.
-VIII-B Federated or On-Premises LLM Deployments
+VII-B Federated or On-Premises LLM Deployments
 Privacy-Focused Approaches
@@
 By conducting model training and inference on-premises or within tightly controlled cloud environments, organizations maintain compliance with confidentiality and data protection regulations while still benefiting from advanced LLM-based QA.
-VIII-C Advanced Multimodal and Integrated SQA
+VII-C Advanced Multimodal and Integrated SQA
 Beyond Text and Code
@@
 By correlating these signals, LLM-driven QA tools could propose deeper, more targeted improvements earlier in the development life cycle.
-VIII-D Standard Updates to Incorporate AI/LLM Methodologies
+VII-D Standard Updates to Incorporate AI/LLM Methodologies
 Call for AI-Specific Guidelines
@@
 Incorporating these mandates into software quality standards will be critical to ensuring both innovation and accountability.
-VIII-E Cross-Industry and Cross-Domain Extensions
+VII-E Cross-Industry and Cross-Domain Extensions
 Knowledge Transfer